Subject: APO 3.5m users committee minutes from 10/5/09 meeting (fwd)

From: Suzanne Hawley

Submitted: Tue, 6 Oct 2009 15:17:11 -0700 (PDT)

Message number: 1148 (previous: 1147, next: 1149 up: Index)

         APO 3.5-m Users Committee Phonecon, 10/05/09

Attending:  Suzanne Hawley, Al Harper, Jon Holtzman, Michael Strauss, Scott 
Anderson, Mark Klaene, John Bally, Russet McMillan, Bill Ketzeback, Bruce 

Absent:  Jon Fulbright, Remy Indebetouw


User feedback, comments from institutional representatives:

Chicago (Al Harper) - Al said he had nothing new to report.

New Mexico State (Jon Holtzman) - Jon reported on a recently received question 
from an NMSU user.  Is is possible to do field-star guiding in horizon mode? 
Russet said it is possible, although there are accuracy limitations, 
particularly for long exposures.  Users are encouraged to try it if they need 
it, although bore-sight guiding and/ or "unguided" exposures may work as well, 
or better.  The Observing specialists will share their experiences with 
field-star horizon-mode guiding amongst themselves so that everyone is up to 

Washington (Scott Anderson) - Scott (and Suzanne) thanked all the site staff 
for their support during the recent visit and orientation of a batch of new UW 

Colorado (John Bally) - A recent CU user of TripleSpec using the long- slit 
mode recommends the need to take darks and off-target sky frames.  This led to 
a discussion of whether the user documentation is explicit on this topic.

===> ACTION:  Russet McMillan will look at the TripleSpec user documentation 
for its adequacy in describing darks and sky calibrations for the long-slit 

Princeton (Michael Strauss) - Michael said that things with the PU users are 
"quiet, happy, and no news."

Johns Hopkins - Suzanne mentioned that she had received a note of appreciation 
from a JHU user who received timely site staff help in unlocking a user 
password blockage.

Virginia - No report.


Discussion of telescope/instruments report:

Mark mentioned that it has only been a short time since the last report.  The 
weather has since been a mixed bag of rain, clouds, clear nights and 
occasionally very good seeing.  The new axes motor controllers have been 
working reasonably well since coming out of shutdown, but there have been 
sporadic dropouts, which the site engineers are working on.  The operations 
building kitchen has received a long-overdue remodeling job.  For the 
instruments, the SPIcam filter wheel is still occasionally hanging up; quick 
modifications are being tried, but the system may need a redesign and overhaul. 
If users notice problems with the SPIcam filter wheel, they should let their 
Observing Specialist know about it when it happens. Also, the GFP developed a 
vacuum leak, which was fixed.  The CIF projects for 2009 are essentially 
completed, and planning is underway for 2010 projects.


3.5-m Telescope, Instruments, and CIF Projects Highlights, 9/10/09 through 

                           Mark Klaene

0) Overview

The first three weeks of observing after the summer shutdown has been a mixed 
blessing.  Weather kept us closed much of the time, but when we were open skies 
were generally photometric and seeing variable but often very good.  Problems 
with the new motors and controllers were much less than expected given the 
minimal testing that was done coming out of shutdown. The kitchen upgrade is 
mostly complete, with new appliances arriving soon.

1) Telescope

There were no unexpected problems this period.  We had one occurrence of the 
altitude axis faulting due to position errors and one due to high wind 
(something not previously tested).  We are addressing these issues during the 
engineering time October 1 and 2.

2) Instruments

The SPIcam filter wheel failed at 120 deg instrument rotation angle. It appears 
that the filter wheel drive circuit is not applying sufficient current during 
the full wheel rotational move. Modifications were made to the circuitry to 
improve this and SPIcam is back in service.  If the problem recurs, a redesign 
of the drive circuitry will be required.

DIS, Echelle, Agile and TripleSpec ran with no problems.

GFP had a vacuum leak and is being pumped on prior to re-cooling later this 
week.  The leak is suspected to be at the valve.

3) CIF projects

With the completion of the CIF projects for 2009, we are currently discussing 
plans for 2010.


2009 Science Highlights (Hawley):

Suzanne said that we have science highlights from Chicago, Washington, 
Colorado, Virginia, and Princeton in hand.  We expect to get highlights from 
NMSU and JHU in the next few days.  Bruce Woodgate has given us a highlight for 
the GFP, and we need to get something for APOLLO.  We will assemble these next 
week for inclusion in the 2009 annual report for the ARC Board.


Time accounting changes (Hawley, McMillan):

Russet said that for several years we have been scheduling the telescope for
usage between sunset and sunrise with a 5-minute cushion, but the usage
accounting in the night logs has been from/to 12 degree twilight.  With several 
instruments now effectively working well into twilight, the usage accounting is 
going to be made consistent in the sense that twilight will be reported in the 
night logs as used (or not) if it could have been effectively used. 
Operationally, there is no change--the twilight time will be accounted for when 
appropriate, and the night log tables will be modified to reflect this.  The 
"Day Use" category will be removed
from the night logs.


ARC futures planning (Hawley):

Suzanne said that discussions are starting about the post-2014 time frame and 
the future of APO.  This will encompass the whole observatory, and we need to 
start reexamining issues pertaining to the 3.5-meter telescope.  A futures 
planning process will be discussed with the ARC Board next month, and we should 
be thinking about topics such as operational modes, institutional alignments, 
new instrumentation, etc.  A futures committee will probably be formed, and it 
will need representation from the 3.5-meter user community.



[open from previous months]:

===> ACTION:  Suzanne needs the 2009 3.5-m telescope science highlights--one 
page of graphics and text per highlight--submitted to Mark Klaene and herself 
not later than 5 October.  Multiple highlights from each institution are 
encouraged, dating back to 1 October '08.

STATUS:  Closed.

[new actions from this meeting]:

===> ACTION:  Russet McMillan will look at the TripleSpec user documentation 
for its adequacy in describing darks and sky calibrations for the long-slit 


Next meeting: The next Users Committee phonecon will be on Monday, 9 November, 
at 8:30 AM Pacific Time.  The agenda and other materials will be sent to the 
committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1148 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to